es we use no. ample size n-30 IS n33o. Use the following table of all 2012 Summer Olympics medalists YOU MAY LEAVE PROBABILIT number 7 please
Answer #1

P(R) = 81/961 = 0.0843 (approx)

P(Gold) = 302/961 = 0.3143 (approx)

P(Russia and Gold) = 24/961 = 0.02497 (approx)

0.0843 * 0.3143 = 0.02688 (approx)

P(Russia and Gold)\tiny \neqP(Russia)*P(Gold)

So, these are not independent events.

Please comment if any doubt. Thank you.